# Can LLMs Generate Tabular Summaries of Science Papers? Rethinking the Evaluation Protocol

### v3

@@ Appendix A Implementation Details
A.1 Prompts Used
We first introduce the prompt used to construct the arXiv2Table benchmark, as explained in Section 4. The main step involves prompting LLM is to collect user demands that describe the purpose of creating the table while remaining contextually self-contained and not revealing the actual schema or values of the table. We use the following prompt to instruct GPT-4o in generating these user demands.
@@ A.1 Prompts Used
Then, for synthesizing QA pairs from a table, we use the following prompt to guide GPT-4o in generating some QA pairs with answers:
@@
A.2 Evaluation Implementations
We access all open-source LLMs via the Hugging Face library Wolf et al. (2020). The models used are meta-llama/Llama-3.3-70B-Instruct, mistralai/Mistral-Large-Instruct-2411, and deepseek-ai/DeepSeek-V3.
For GPT models, we access them via the official OpenAI Batch API [^2] https://platform.openai.com/docs/guides/batch. The models used are gpt-4o-mini-2024-07-18 and gpt-4o-2024-08-06.
Note that the DeepSeek model family has a context window limit of 64K tokens, whereas the others have a limit of 128K tokens. The generation temperature is set to 0.5 for all experiments. All experiments are repeated twice and the average performance is reported.
Statistic Paper Count Column Count Distractor Count
Min 1 2 4
Max 35 13 10
Mean 3.65 3.56 5.21
Total 7158 6967 10196
Figure 7: Distribution of number of papers in each table.
A.3 Computational Cost Comparisons
Method Success Rate #Tokens Avg. Runtime
Baseline 1 48.2% 128K 37
Baseline 2 98.2% 167K 118
Newman et al. 99.7% 110K 208
Ours 100.0% 118K 194
Table 6: Computational cost and efficiency metrics across different generation methods using LLaMA-3.3. We report the generation success rate, average token usage, and average runtime (s) per table.
Appendix B Annotation Details
To ensure the high quality of our human annotations, we implement strict quality control measures. First, we select only postgraduate students with research experience in computer science to ensure they are familiar with relevant topics. All selected annotators undergo qualification rounds, and we invite only those who demonstrate satisfactory performance to serve as our main annotators.
For each task, we provide workers with comprehensive task explanations in layman’s terms to enhance their understanding. Additionally, we offer detailed definitions and multiple examples for each choice to help annotators make informed decisions. Each entry requires the worker to provide a binary vote on whether the paper should be excluded or not. Our annotation interface is shown in Figure 8.
To ensure comprehension, we require annotators to confirm that they have thoroughly read the instructions by ticking a checkbox before starting the annotation task. We also manually monitor the performance of annotators throughout the annotation process and provide feedback based on common errors. Spammers or underperforming workers are disqualified. As described in Section 4.2, the inter-annotator agreement supports the quality of our collected annotations.
@@ Appendix B Annotation Details
Appendix C Case Studies
Table 7 presents randomly sampled examples of original table captions alongside their improved user demands, demonstrating how refining vague captions enhances specificity and ensures more structured table generation. The findings highlight that well-defined user demands help capture key aspects of table construction, leading to more informative and targeted tabular representations.
Table 8 illustrates schema, unary value, and pairwise value questions designed to assess the quality of generated tables, ensuring alignment with ground-truth information. The results reveal that this QA-based evaluation effectively quantifies schema retention, individual value accuracy, and consistency in relationships, providing a structured approach for benchmarking table generation models.
In addition, we present two pairs of ground-truth and generated tables as examples for a case study on table generation, as shown in Table 9. From these tables, we observe that the generation process is capable of incorporating many useful columns, thereby enriching the available information. For instance, in the first example, the generated table introduces new columns such as Number of Images, Number of Subjects, and Avg. Images per Subject, which add valuable quantitative insights beyond the original ground truth table. However, it is also evident that some columns present in the ground truth, like the Evaluation Metric, are not fully covered in the generated version. In the second example, the user demand for detailed descriptions has led to a generated table with numerous specific columns, including ID, Method Used, Performance Metric, and Results Achieved. Although these additional details enhance the descriptive quality of the table, they also suggest a potential issue: the need for further polishing and refinement of the user demand to balance detail with clarity.
@@ Appendix C Case Studies
Comparison of Trajectory and Path Planing Approach Generate a table that compares different trajectory and path planning approaches, focusing on their collision avoidance techniques, benefits, limitations, and applicable scenarios. The table should include detailed columns to capture these aspects for each method mentioned in the relevant papers.
@@
Publications with deep-learning focused sampling methods. We cluster the papers based on the space the sample through and how the samples are evaluated. Some approaches further consider an optional refinement stage. Create a table that categorizes publications focused on deep-learning-based sampling methods for grasp detection, organizing them by the space in which samples are generated, the evaluation criteria used, and whether a refinement stage is included. The table should provide a comprehensive yet concise overview of the methodological variations and enhancements across different papers.
Categorization of textual explanation methods. Create a table that categorizes the methods used for providing textual explanations in visual question answering systems, focusing on the types of texts generated and the reasoning processes employed. The table should use succinct columns to differentiate between these methodological aspects for each paper.
Metadata of the three benchmarks that we focus on. XSumSota is a combined benchmark of cite:1400aac and cite:d420ef8 for summaries generated by the state-of-the-art summarization models. Create a table that details the metadata for three summarization benchmarks, focusing on the composition of annotators, the dataset sizes for validation and testing, and the distribution of positive and negative evaluations. The table should provide a comprehensive comparison across these aspects for each benchmark.
Review of open access ground-based forest datasets Create a table that reviews various open-access forest datasets, focusing on the publication and data recording years, types of data collected, and their applicability to specific forestry-related tasks. The table should offer a concise summary of each dataset’s attributes, including the number of classification categories and geographical location.
Comparison of existing consistency-type models. Create a table that compares different models focusing on their purpose, the trajectory they follow, the main objects they equate, and their methodological approach. The table should provide detailed insights into how each model addresses consistency issues, drawing from specified papers.
Table 7: Randomly sampled examples of the original captions and their corresponding improved user demands. Most captions are relatively short and may be vague without the full table’s content.
Schema Unary Value Pairwise Value
Is Dataset included in the table schema? Is CL, TL the loss function for paper CN-LexNet? Is ResNet-v2 using more evaluation metrics than GAN?
Is Model Architecture included in the table schema? Is GPT-4o the model used for multimodal understanding? Does GPT-4o have a larger parameter size than LLaMA-2?
Is Training Dataset included in the table schema? Is ImageNet the dataset used for training ResNet? Is ResNet trained on more samples than EfficientNet?
Is Performance Metric included in the table schema? Is BLEU-4 the evaluation metric for MT-BERT? Does BERT outperform LSTM on BLEU-4 score?
Is Activation Function included in the table schema? Is ReLU the activation function used in Transformer? Is GELU smoother than ReLU in function continuity?
Is Optimization Algorithm included in the table schema? Is Adam the optimizer used for training BERT? Does Adam converge faster than SGD for BERT training?
Is Pretraining Task included in the table schema? Is Masked Language Modeling the pretraining task for BERT? Does BERT use a more complex pretraining strategy than GPT?
Is Hyperparameter included in the table schema? Is the learning rate set to 0.001 for training ViT? Does ViT use a higher learning rate than ResNet?
Is Hardware Accelerator included in the table schema? Is TPU used for training T5? Do TPUs provide faster training than GPUs for T5?
Table 8: Randomly sampled examples of schema, unary value, and pairwise value questions used to evaluate the quality of generated tables. Each row contains three related questions derived from the same table.
@@
208138178 Adaptive dual encoding framework VisDial Visual dialogue State-of-the-art results Generative 2023

### v5

@@ Appendix B Implementation Details
B.1 Prompts Used
We provide all prompts used in our benchmark construction and evaluation. In benchmark construction, the key steps are (i) rewriting captions into user demands without leaking schema/values and (ii) synthesizing QA pairs from the ground-truth table. In evaluation, we normalize tables, canonicalize headers, decouple QA synthesis from answering, and run a reverse-QA pass for precision.
User demand rewrite (benchmark construction). This prompt rewrites terse/arXiv-style captions into contextually self-contained user demands that specify the table’s purpose without leaking column names or specific values. It improves realism by emulating well-specified but abstract requests while keeping evaluation fair.
@@ B.1 Prompts Used
Leak check for user demands (guardrail). Immediately after rewriting, we run a leak check to ensure the demand does not expose schema labels or table values, and to auto-rewrite if needed. This keeps construction oracle-free and reproducible.
You are given: (i) a user demand written from a caption, and (ii) the target table (schema and a few cell values).
Decide if the user demand leaks any column names or specific cell values, or directly paraphrases them.
Return {ACCEPT, REWRITE} and a one-sentence reason. If REWRITE, produce a version that removes leaked tokens while staying specific and self-contained.
Output JSON only: {"decision": "...", "reason": "...", "demand": "..." }
QA synthesis from ground truth (recall side). This prompt generates binary QA pairs from the gold table to test whether a system table retains schema, values, and pairwise relations. It provides full coverage for schema/unary and a sampled, diverse set for pairwise.
@@
Answering gold QAs on the system table (decoupled answerer). When we decouple synthesis from answering (for evaluator-independence checks and robustness), we use an answer-only prompt that strictly binds answers to the system table content.
Input: (i) a normalized generated table, and (ii) a list of binary Yes/No QAs synthesized from the ground-truth table.
Answer each QA using only the generated table. If explicitly supported and consistent, answer "yes"; otherwise "no".
Return one lowercase token per line: yes/no (no explanations).
Reverse-QA synthesis from the system table (precision side). To measure precision, we synthesize QAs from the system table and answer them on the gold table. This credits only content supported by the gold table and rejects unsupported “novel” entries.
Given a normalized generated table, create binary Yes/No QAs for: (i) schema presence, (ii) specific cell values, and (iii) pairwise relations (10 pairs).
Write unambiguous questions that can be answered solely from this generated table. Avoid trivial or duplicate QAs.
Return JSON list: ["type":"schema|unary|pairwise","q":"..." ...]
Answering reverse-QAs on the gold table (precision side). This answer-only prompt binds answers to the gold table. By instruction, any system-only content is answered “no”.
Input: (i) a normalized gold table, and (ii) a list of Yes/No QAs synthesized from the generated table.
Answer each QA using only the gold table. Novel content not present in gold must be answered "no".
Return one lowercase token per line: yes/no.
QA validation and deduplication (quality filter). Before answering, we filter QAs to remove ill-formed, non-binary, trivial, or duplicate items. This reduces evaluator brittleness and ensures consistent scoring.
Given a list of binary QAs, remove any that are ill-formed, non-binary, trivially true/false, or duplicates.
Return the filtered list as JSON in the same schema and include a "removed" list with reasons.
Table normalization (robust parsing). We normalize both gold and system tables to a rectangular CSV with consistent headers and “N/A” for missing entries so that downstream prompts and scripts operate deterministically.
You are given a table in arbitrary Markdown/LaTeX/CSV. Normalize it into a rectangular CSV with a header row, one row per paper.
-- Trim whitespace; collapse multi-line cells; preserve units and text verbatim; fill missing cells with "N/A".
-- Do not infer new columns or values.
Return CSV only (no commentary).
Schema alias canonicalization (header mapping). We canonicalize header synonyms (e.g., “Dataset” vs “Data”) before computing schema scores to reduce false mismatches due to phrasing.
Given two header lists (gold vs system), produce a one-to-one or one-to-many mapping of system headers to gold headers when they are semantically equivalent.
Rules: prefer exact matches; allow synonyms; keep units consistent; do not map if semantics differ.
Return JSON: ["system": "...", "gold": "...", "justification": "..." ...]
@@
B.2 Evaluation Implementations
We access all open-source LLMs via the Hugging Face library Wolf et al. (2020). The models used are meta-llama/Llama-3.3-70B-Instruct, mistralai/Mistral-Large-Instruct-2411, and deepseek-ai/DeepSeek-V3. For GPT models, we access them via the official OpenAI Batch API [^3] https://platform.openai.com/docs/guides/batch. The models used are gpt-4o-mini-2024-07-18 and gpt-4o-2024-08-06. Note that the DeepSeek model family has a context window limit of 64K tokens, whereas the others have a limit of 128K tokens. The generation temperature is set to 0.5 for all experiments. All experiments are repeated twice and the average performance is reported.
We normalize both gold and generated tables to CSV grids with a single header row and “N/A” for missing cells, then apply schema-alias canonicalization before scoring. For the recall side, we synthesize QAs from the gold table and answer them on the system table; for the precision side, we synthesize QAs from the system table and answer them on the gold table. Answers are strictly “yes”/“no”; we lowercase, strip punctuation, and parse the first token if extra text is emitted. Schema and unary QAs are exhaustive; for pairwise, we sample 10 pairs per table while avoiding duplicates and promoting coverage over distinct columns. Randomization uses fixed seeds when supported; otherwise we repeat evaluation twice and average.
For cross-evaluator checks, we decouple QA synthesis and answering and swap the evaluator models as reported in Appendix A.3. Decoding parameters use temperature 0.5, top_p 1.0, no repetition penalties, and a max_new_tokens budget that comfortably covers the QA list; timeouts are 120s per call with up to two retries on transient errors. Significance testing for main-table comparisons uses table-level bootstrap (10,000 resamples) with 95% confidence intervals; full CIs and prompt JSON I/O schemas will be released with code. For the additional cross-family evaluator swap, we also use Qwen3-30B-A3B-Instruct under the same decoding settings Wang et al. (2026, 2024a, 2025b, 2023b, 2023a, 2025a); Wang and Song (2025).
Appendix C Annotation Details
To ensure high-quality annotations, we recruited 19 trained graduate-level annotators with computer-science research experience and admitted them only after passing qualification rounds. For each candidate paper, annotators received clear, layman-friendly instructions with definitions and multiple examples, then confirmed they had read them via a checkbox before starting. Using the interface in Figure 8, each instance was judged with a binary {include, exclude} decision relative to the user demand and the known reference papers (title+abstract basis). Every instance received two independent labels; on disagreement, two additional adjudicators resolved to consensus. We continuously monitored performance, provided targeted feedback on common errors, and removed spammers or underperformers. The study ran for eleven days; first-pass agreement reached 94% pairwise with Fleiss’ $\kappa{=}0.73$, the self-reported median time per decision was about seven minutes, and annotators were compensated above the local minimum wage. The final corpus contains 10,196 curated distractor labels across 1,957 tables (with 10 initial candidates per table before filtering), supporting the quality reported in Section 4.2.
@@ Appendix C Annotation Details
Appendix D Case Studies
Table 16 shows examples of original captions and rewritten user demands, illustrating how short or context-dependent captions can be transformed into self-contained requests that better specify the intended comparison goal. Table 17 provides illustrative examples of schema, unary, and pairwise questions used by our utilization-based evaluation to measure schema coverage, factual retention, and relational consistency.
We additionally include two example pairs of ground-truth and generated tables in Table 18. These examples highlight common behaviors in literature-review table generation: systems can enrich tables by adding helpful organizing fields, but may also omit important attributes from the ground truth or introduce overly fine-grained fields that dilute the central comparison. Overall, the examples underscore the importance of jointly handling paper selection, schema induction, and value grounding.
@@ Appendix D Case Studies
Comparison of trajectory and path planning approaches I need a table that compares recent trajectory/path planning methods, emphasizing how they handle safety constraints (e.g., collision avoidance) and what assumptions they make about the environment. Please summarize the key strengths, limitations, and typical application settings for each approach.
@@
Publications with deep-learning focused sampling methods. We cluster the papers based on the space the sample through and how the samples are evaluated. Some approaches further consider an optional refinement stage. Please create a table that organizes learning-based sampling methods by how candidates are generated and how they are scored, so it is easy to see the main design choices across papers. If a method uses an extra refinement step or additional supervision, highlight that in the comparison.
Categorization of textual explanation methods. I want a table that groups approaches for producing textual explanations by what kind of explanation they generate and what evidence they rely on (e.g., rationales, templates, retrieved facts). The goal is to quickly compare how different families of methods justify their predictions.
Metadata of the three benchmarks that we focus on. XSumSota is a combined benchmark of cite:1400aac and cite:d420ef8 for summaries generated by the state-of-the-art summarization models. Please produce a table that contrasts these benchmarks in terms of what they measure and how they are evaluated, including how labels are obtained and what the evaluation protocol looks like. I want to understand which benchmark is most suitable for comparing modern summarization systems.
Review of open access ground-based forest datasets I need a table summarizing open-access forest datasets, focusing on what is recorded, when and where data are collected, and what tasks each dataset supports. The table should make it easy to choose a dataset for a specific forest monitoring objective.
Comparison of existing consistency-type models. Please construct a table comparing consistency-oriented models by what notion of consistency they enforce and how they operationalize it (objective, constraints, or training signal). The table should make clear how each model differs in assumptions and what settings it is intended for.
Table 16: Examples of original captions and rewritten user demands. User demands are written to be self-contained and goal-oriented while avoiding direct leakage of gold headers or specific cell values.
Schema QA (presence of a field) Unary QA (a specific entry) Pairwise QA (a relation)
Is the evaluation benchmark included in the table schema? Is ZsRE listed as an evaluation benchmark for ROME? Does ROME achieve higher reported accuracy than MEND on ZsRE?
Is the training signal/objective described in the table schema? Is consistency regularization listed as a training signal for at least one method in the table? Do methods that use consistency regularization report better performance than those that do not on the same benchmark?
Is the data source or dataset type included in the table schema? Is ImageNet listed as a dataset used in any compared method? Is Method A evaluated on more datasets than Method B in the table?
[…]
Is the publication year or timeframe included in the table schema? Is 2023 listed as the publication year for Method X? Are post-2022 methods reported as outperforming pre-2020 methods on the same metric in the table?
Table 17: Illustrative examples of schema/unary/pairwise questions. In the benchmark, questions are synthesized from each ground-truth table and are answerable by reading the corresponding comparison table.
@@
208138178 Adaptive dual encoding framework VisDial Visual dialogue Accuracy State-of-the-art results Generative